OpenStudy (anonymous):

WHAT IS 170 PERCENT CONVERTED INTO A FRACTION?

OpenStudy (anonymous):

percent means out of 100 170 out of 100 ex. 1 out of 2 = 1/2 so 170 out of 100 =? then simplify

OpenStudy (anonymous):

1.7

OpenStudy (anonymous):

170/100 = your fraction which should come to 17/10